If you needed to perform a reaction in a controlled pH environment that was fairly basic,...

  1. If you needed to perform a reaction in a controlled pH environment that was fairly basic, you might choose to use the ammonium and ammonia buffer system. How many grams of solid ammonium chloride would you have to add to 250 liters of 0.165M NH3 to obtain a buffered solution of pH = 8.85?

(Kb of NH3 = 1.8 x 10-5)
Answer #1

Given :- Concentration of NH₃ = 0.165m Volume of NH3 = 250 litres pH of buffer solution 58-85 Kb of NH₃ = 118010 1.8 X/ SolutpOH = 14- pH pOH = 14 - 8.85 [pOH= 5:15] Step (li) calculate pro рк, -а, К. - log (1.8x105) = -log 1:8 + 5logio -0.2553 + 5 4We get, J: 2.5704 {NH3 This gives the ratio between the concentration of NHAT and concentration of NH3 [NH4+] = 2.5704 . [NHdoes not change after the number of moles solution is calculated Moles n= addition of NHA of NHcl in the using the formula, W

5671g of solid ammonium chloride add to 250 liters of 0.165M NH3 to obtain a buffered solution of pH = 8.85